OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = 3, point (2, 1)

OpenStudy (anonymous):

given the gradient m , and a point (x1 , y1) the equation of the line is: y- y1 = m(x -x1)

OpenStudy (anonymous):

wat is the question?

OpenStudy (anonymous):

y-1 = 3(x -2)

OpenStudy (anonymous):

Y=3X-5?

OpenStudy (anonymous):

Using slope intercept form, y= mx + b and the fact that m = 3, we get y = 3x + b. Now use the specific point you're given to solve for b. Plugging in (2, 1) gives: 1 = 3(2) +b b = 1-6 = -5 so m = 3, b = -5 and the equation of the line is y = 3x - 5

OpenStudy (anonymous):

Write the equation of the line given the slope and the y-intercept. Type your answer in the box without skipping any spaces. m = 2, y-intercept (0, -9) I know this is the same thing but becuse it has a zero how would u go about solving this?

OpenStudy (anonymous):

The y intercept is the point where the line crosses the y axis. Because it crosses the y axis, that means the x coordinate of that point will always be 0. The y coordinate of the point is the intercept, or the b in "y = mx + b"

OpenStudy (anonymous):